# Agentic Context Engineering: Evolving Contexts for Self-Improving Language Models

### v2

@@ Abstract
Large language model (LLM) applications such as agents and domain-specific reasoning increasingly rely on context adaptation—modifying inputs with instructions, strategies, or evidence, rather than weight updates. Prior approaches improve usability but often suffer from brevity bias, which drops domain insights for concise summaries, and from context collapse, where iterative rewriting erodes details over time. Building on the adaptive memory introduced by Dynamic Cheatsheet, we introduce ACE (Agentic Context Engineering), a framework that treats contexts as evolving playbooks that accumulate, refine, and organize strategies through a modular process of generation, reflection, and curation. ACE prevents collapse with structured, incremental updates that preserve detailed knowledge and scale with long-context models. Across agent and domain-specific benchmarks, ACE optimizes contexts both offline (e.g., system prompts) and online (e.g., agent memory), consistently outperforming strong baselines: +10.6% on agents and +8.6% on finance, while significantly reducing adaptation latency and rollout cost. Notably, ACE could adapt effectively without labeled supervision and instead by leveraging natural execution feedback. On the AppWorld leaderboard, ACE matches the top-ranked production-level agent on the overall average and surpasses it on the harder test-challenge split, despite using a smaller open-source model. These results show that comprehensive, evolving contexts enable scalable, efficient, and self-improving LLM systems with low overhead.
@@ 1 Introduction
Figure 1: Overall Performance Results. Our proposed framework, ACE, consistently outperforms strong baselines across agent and domain-specific reasoning tasks.
@@
Modern AI applications based on large language models (LLMs), such as LLM agents [yao2023react, yang2024swe] and compound AI systems [zaharia2024compoundGS], increasingly depend on context adaptation. Instead of modifying model weights, context adaptation improves performance after model training by incorporating clarified instructions, structured reasoning steps, or domain-specific input formats directly into the model’s inputs. Contexts underpin many AI system components, including system prompts that guide downstream tasks [opsahl2024optimizing, agrawal2025gepa], memory that carries past facts and experiences [suzgun2025dynamic, xu2025mem], and factual evidence that reduces hallucination and supplements knowledge [asai2024self].
@@
Adapting through contexts rather than weights offers several key advantages. Contexts are interpretable and explainable for users and developers [wei2022chain, wang2022self], allow rapid integration of new knowledge at runtime [lewis2020retrieval, borgeaud2022improving], and can be shared across models or modules in a compound system [khot2022decomposed]. Meanwhile, advances in long-context LLMs [peng2023yarn] and context-efficient inference such as KV cache reuse [gim2024prompt, yao2025cacheblend] are making context-based approaches increasingly practical for deployment. As a result, context adaptation is emerging as a central paradigm for building capable, scalable, and self-improving AI systems.
@@
Despite this progress, existing approaches to context adaptation face two key limitations. First, a brevity bias: many prompt optimizers prioritize concise, broadly applicable instructions over comprehensive accumulation. For example, GEPA [agrawal2025gepa] highlights brevity as a strength, but such abstraction can omit domain-specific heuristics, tool-use guidelines, or common failure modes that matter in practice [gao2025prompt]. This objective aligns with validation metrics in some settings, but often fails to capture the detailed strategies required by agents and knowledge-intensive applications. Second, context collapse: methods that rely on monolithic rewriting by an LLM often degrade into shorter, less informative summaries over time, causing sharp performance declines (Figure 2). In domains such as interactive agents [trivedi2024appworld, patil2024gorilla, zhang2024caravan], domain-specific programming [ye2023generating, zhang2025adaptive], and financial or legal analysis [loukas2022finer, guha2023legalbench, wang2025finlora], strong performance depends on retaining detailed, task-specific knowledge rather than compressing it away.
@@
As applications such as agents and knowledge-intensive reasoning demand greater reliability, recent work has shifted toward saturating contexts with abundant, potentially useful information [jiang2025putting, chung2025long, chen2025flora], enabled by advances in long-context LLMs [peng2023yarn, mao2024lift]. We argue that contexts should function not as concise summaries, but as comprehensive, evolving playbooks—detailed, inclusive, and rich with domain insights. Unlike humans, who often benefit from concise generalization, LLMs are more effective when provided with long, detailed contexts and can distill relevance autonomously [jiang2025putting, liu2025selfelicit, suzgun2025dynamic]. Thus, instead of compressing away domain-specific heuristics and tactics, contexts should preserve them, allowing the model to decide what matters at inference time.
@@
To address these limitations, we introduce ACE (Agentic Context Engineering), a framework for comprehensive context adaptation in both offline settings (e.g., system prompt optimization) and online settings (e.g., test-time memory adaptation). Rather than compressing contexts into distilled summaries, ACE treats them as evolving playbooks that accumulate and organize strategies over time. Building on the agentic architecture of Dynamic Cheatsheet [suzgun2025dynamic], ACE incorporates a modular workflow of generation, reflection, and curation, while adding structured, incremental updates guided by a grow-and-refine principle. This design preserves detailed, domain-specific knowledge, prevents context collapse, and yields contexts that remain comprehensive and scalable throughout adaptation.
@@
We evaluate ACE on two categories of LLM applications that most benefit from comprehensive, evolving contexts: (1) agents [trivedi2024appworld], which require multi-turn reasoning, tool use, and environment interaction, where accumulated strategies can be reused across episodes; and (2) domain-specific benchmarks, which demand specialized tactics and knowledge, where we focus on financial analysis [loukas2022finer, wang2025finlora]. Our key findings are:
@@
    ACE is able to construct effective contexts without labeled supervision, instead leveraging execution feedback and environment signals—key ingredients for self-improving LLMs and agents.
@@
    On the AppWorld benchmark leaderboard [AppWorldLeaderboard], ACE matches the top-ranked production-level agent IBM-CUGA [marreed2025towards] (powered by GPT-4.1) on average and surpasses it on the harder test-challenge split, while using a smaller open-source model (DeepSeek-V3.1).
@@
    ACE requires significantly fewer rollouts and lower dollar costs, and achieves 86.9% lower adaptation latency (on average) than existing adaptive methods, demonstrating that scalable self-improvement can be achieved with both higher accuracy and lower overhead.
@@ 2.1 Context Adaptation
Context adaptation (or context engineering) refers to methods that improve model behavior by constructing or modifying inputs to an LLM, rather than altering its weights. The current state of the art leverages natural language feedback [shinn2023reflexion, yuksekgonul2024textgrad, agrawal2025gepa]. In this paradigm, a language model inspects the current context along with signals such as execution traces, reasoning steps, or validation results, and generates natural language feedback on how the context should be revised. This feedback is then incorporated into the context, enabling iterative adaptation. Representative methods include Reflexion [shinn2023reflexion], which reflects on failures to improve agent planning; TextGrad [yuksekgonul2024textgrad], which optimizes prompts via gradient-like textual feedback; GEPA [agrawal2025gepa], which refines prompts iteratively based on execution traces and achieves strong performance, even surpassing reinforcement learning approaches in some settings; and Dynamic Cheatsheet [suzgun2025dynamic], which constructs an external memory that accumulates strategies and lessons from past successes and failures during inference. These natural language feedback methods represent a major advance, offering flexible and interpretable signals for improving LLM systems beyond weight updates.
@@ 2.2 Limitations of Existing Context Adaptation Methods
The Brevity Bias.
A recurring limitation of context adaptation methods is brevity bias: the tendency of optimization to collapse toward short, generic prompts. Gao et al. [gao2025prompt] document this effect in prompt optimization for test generation, where iterative methods repeatedly produced near-identical instructions (e.g., "Create unit tests to ensure methods behave as expected"), sacrificing diversity and omitting domain-specific detail. This convergence not only narrows the search space but also propagates recurring errors across iterations, since optimized prompts often inherit the same faults as their seeds. More broadly, such bias undermines performance in domains that demand detailed, context-rich guidance—such as multi-step agents, program synthesis, or knowledge-intensive reasoning—where success hinges on accumulating rather than compressing task-specific insights.
@@
Context Collapse.
In a case study on the AppWorld benchmark [trivedi2024appworld], we observe a phenomenon we call context collapse, which arises when an LLM is tasked with fully rewriting the accumulated context at each adaptation step. As the context grows large, the model tends to compress it into much shorter, less informative summaries, causing a dramatic loss of information. For instance, at step 60 the context contained 18,282 tokens and achieved an accuracy of 66.7, but at the very next step it collapsed to just 122 tokens, with accuracy dropping to 57.1—worse than the baseline accuracy of 63.7 without adaptation. While we highlight this through Dynamic Cheatsheet [suzgun2025dynamic], the issue is not specific to that method; rather, it reflects a fundamental risk of end-to-end context rewriting with LLMs, where accumulated knowledge can be abruptly erased instead of preserved.
@@ 3 Agentic Context Engineering (ACE)
We present ACE (Agentic Context Engineering), a framework for scalable and efficient context adaptation in both offline (e.g., system prompt optimization) and online (e.g., test-time memory adaptation) scenarios. Instead of condensing knowledge into terse summaries or static instructions, ACE treats contexts as evolving playbooks that continuously accumulate, refine, and organize strategies over time. Building on the agentic design of Dynamic Cheatsheet [suzgun2025dynamic], ACE introduces a structured division of labor across three roles (Figure 4): the Generator, which produces reasoning trajectories; the Reflector, which distills concrete insights from successes and errors; and the Curator, which integrates these insights into structured context updates. This mirrors how humans learn—experimenting, reflecting, and consolidating—while avoiding the bottleneck of overloading a single model with all responsibilities.
To address the limitations of prior methods discussed in §2.2—notably brevity bias and context collapse—ACE introduces three key innovations: (1) a dedicated Reflector that separates evaluation and insight extraction from curation, improving context quality and downstream performance (§4.5); (2) incremental delta updates3.1) that replace costly monolithic rewrites with localized edits, reducing both latency and compute cost (§4.6); and (3) a grow-and-refine mechanism (§3.2) that balances steady context expansion with redundancy control.
@@ 3.1 Incremental Delta Updates
A core design principle of ACE is to represent context as a collection of structured, itemized bullets, rather than a single monolithic prompt. The concept of a bullet is similar to the concept of a memory entry in LLM memory frameworks like Dynamic Cheatsheet [suzgun2025dynamic] and A-MEM [xu2025mem], but builds on top of that and consists of (1) metadata, including a unique identifier and counters tracking how often it was marked helpful or harmful; and (2) content, capturing a small unit such as a reusable strategy, domain concept, or common failure mode. When solving new problems, the Generator highlights which bullets were useful or misleading, providing feedback that guides the Reflector in proposing corrective updates.
This itemized design enables three key properties: (1) localization, so only the relevant bullets are updated; (2) fine-grained retrieval, so the Generator can focus on the most pertinent knowledge; and (3) incremental adaptation, allowing efficient merging, pruning, and de-duplication during inference.
@@ 4 Results
    Enabling High-Performance, Self-Improving Agents. ACE enables agents to self-improve by dynamically refining their input context. It boosts accuracy on the AppWorld benchmark by up to 17.1% by learning to engineer better contexts from execution feedback alone, without needing ground-truth labels. This context-driven improvement allows a smaller, open-source model to match the performance of the top-ranked proprietary agent on the leaderboard. (§4.3)
@@
    Effective by Design. Ablation studies confirm our design choices are key to success, with components like the Reflector and multi-epoch refinement each contributing substantial performance gains. (§4.5)
@@ 4.1 Tasks and Datasets
We evaluate ACE on two categories of LLM applications that benefit most from a comprehensive and evolving context: (1) agent benchmarks, which require multi-turn reasoning, tool use, and environment interaction, where agents can accumulate and reuse strategies across episodes and environments; and (2) domain-specific benchmarks, which demand mastery of specialized concepts and tactics, where we focus on financial analysis as a case study.
@@
    LLM Agent: AppWorld [trivedi2024appworld] is a suite of autonomous agent tasks involving API understanding, code generation, and environment interaction. It provides a realistic execution environment with common applications and APIs (e.g., email, file system) and tasks of two difficulty levels (normal and challenge). A public leaderboard [AppWorldLeaderboard] tracks performance, where, at the time of submission, the best system achieved only 60.3% average accuracy, highlighting the benchmark’s difficulty and realism.
@@
    Financial Analysis: FiNER [loukas2022finer] and Formula [wang2025finlora] test LLMs on financial reasoning tasks that rely on the eXtensible Business Reporting Language (XBRL). FiNER requires labeling tokens in XBRL financial documents with one of 139 fine-grained entity types, a key step for financial information extraction in regulated domains. Formula focuses on extracting values from structured XBRL filings and performing computations to answer financial queries, i.e., numerical reasoning.
Evaluation Metrics.
For AppWorld, we follow the official benchmark protocol and report Task Goal Completion (TGC) and Scenario Goal Completion (SGC) on both the test-normal and test-challenge splits. For FiNER and Formula, we follow the original setup and report accuracy, measured as the proportion of predicted answers that exactly match the ground truth.

### v3

@@ Abstract
Large language model (LLM) applications such as agents and domain-specific reasoning increasingly rely on context adaptation: modifying inputs with instructions, strategies, or evidence, rather than weight updates. Prior approaches improve usability but often suffer from brevity bias, which drops domain insights for concise summaries, and from context collapse, where iterative rewriting erodes details over time. We introduce ACE (Agentic Context Engineering), a framework that treats contexts as evolving playbooks that accumulate, refine, and organize strategies through a modular process of generation, reflection, and curation. ACE prevents collapse with structured, incremental updates that preserve detailed knowledge and scale with long-context models. Across agent and domain-specific benchmarks, ACE optimizes contexts both offline (e.g., system prompts) and online (e.g., agent memory), consistently outperforming strong baselines: +10.6% on agents and +8.6% on finance, while significantly reducing adaptation latency and rollout cost. Notably, ACE could adapt effectively without labeled supervision and instead by leveraging natural execution feedback. On the AppWorld leaderboard, ACE matches the top-ranked production-level agent on the overall average and surpasses it on the harder test-challenge split, despite using a smaller open-source model. These results show that comprehensive, evolving contexts enable scalable, efficient, and self-improving LLM systems with low overhead.
@@ 1 Introduction
Figure 1: Overall Performance Results. Our proposed framework, ACE, consistently outperforms strong baselines across agent and domain-specific tasks.
@@
Modern AI applications based on large language models (LLMs), such as LLM agents (yao2023react; yang2024swe) and compound AI systems (zaharia2024compoundGS), increasingly depend on context adaptation. Instead of modifying model weights, context adaptation improves performance after model training by incorporating clarified instructions, structured reasoning steps, or domain-specific input formats directly into the model’s inputs. Contexts underpin many AI system components, including system prompts that guide downstream tasks (opsahl2024optimizing; agrawal2025gepa), memory that carries past facts and experiences (suzgun2025dynamic; xu2025mem), and factual evidence that reduces hallucination and supplements knowledge (asai2024self).
@@
Adapting through contexts rather than weights offers several key advantages. Contexts are interpretable and explainable for users and developers (wei2022chain; wang2022self), allow rapid integration of new knowledge at runtime (lewis2020retrieval; borgeaud2022improving), and can be shared across models or modules in a compound system (khot2022decomposed). Meanwhile, advances in long-context LLMs (peng2023yarn) and context-efficient inference such as KV cache reuse (gim2024prompt; yao2025cacheblend) are making context-based approaches increasingly practical for deployment. As a result, context adaptation is emerging as a central paradigm for building capable, scalable, and self-improving AI systems.
@@
Despite this progress, existing approaches to context adaptation face two limitations. First, brevity bias: many prompt optimizers prioritize concise applicable instructions over comprehensive accumulation. For example, GEPA (agrawal2025gepa) highlights brevity as a strength, but such abstraction can omit domain-specific heuristics, tool-use guidelines, or common failure modes that matter in practice (gao2025prompt). This objective aligns with validation metrics in some settings, but often fails to capture the detailed strategies required by agents and knowledge-intensive applications. Second, context collapse: methods that rely on monolithic rewriting by an LLM often degrade into shorter, less informative summaries over time, causing sharp performance declines (Figure 2). In domains such as interactive agents (trivedi2024appworld; patil2024gorilla; zhang2024caravan), domain-specific programming (ye2023generating; zhang2025adaptive; zhang2025accelopt; mang2025frontiercs), and financial or legal analysis (loukas2022finer; guha2023legalbench; wang2025finlora), strong performance depends on retaining detailed, task-specific knowledge rather than compressing it away.
@@
As applications like agents and knowledge-intensive reasoning demand greater reliability, recent work has shifted toward saturating contexts with abundant, potentially useful information (jiang2025putting; chung2025long; chen2025flora), enabled by advances in long-context LLMs (peng2023yarn; mao2024lift). We argue that contexts should function not as concise summaries, but as comprehensive, structured playbooks that are detailed, inclusive, and rich with domain insights. Unlike humans, who often benefit from concise generalization, LLMs are more effective when provided with long, detailed contexts and can distill relevance autonomously (jiang2025putting; liu2025selfelicit; suzgun2025dynamic). Thus, instead of compressing away domain-specific heuristics and tactics, contexts should preserve them, allowing the model to decide what matters during inference time.
@@
To address these limitations, we introduce ACE (Agentic Context Engineering), a framework for comprehensive context adaptation in both offline settings (e.g., system prompt optimization) and online settings (e.g., test-time memory adaptation). Rather than compressing contexts into distilled summaries, ACE treats them as evolving playbooks that accumulate and organize strategies over time. By design, ACE incorporates a modular workflow of generation, reflection, and curation, while adding structured, incremental updates guided by a grow-and-refine principle. This design preserves detailed, domain-specific knowledge, prevents context collapse, and yields contexts that remain comprehensive and scalable throughout adaptation.
@@
We evaluate ACE on two categories of LLM applications that most benefit from comprehensive, evolving contexts: (1) agents (trivedi2024appworld), which require multi-turn reasoning, tool use, and environment interaction, where accumulated strategies can be reused across episodes; and (2) domain-specific benchmarks, which demand specialized tactics and knowledge, like financial analysis (loukas2022finer; wang2025finlora). Our key findings are:
@@
    ACE is able to construct effective contexts without labeled supervision, instead leveraging execution feedback and environment signals, key ingredients for self-improving LLMs and agents.
@@
    On the AppWorld benchmark leaderboard (AppWorldLeaderboard), ACE surpasses the top-1-ranked production-level agent IBM-CUGA (marreed2025towards) (powered by GPT-4.1) while using an open-source model (DeepSeek-V3.1).
@@
    ACE requires significantly fewer rollouts and achieves lower adaptation latency than existing adaptive methods, demonstrating that scalable self-improvement can be achieved with both higher accuracy and lower cost.
@@ 2.1 Context Adaptation
Context adaptation (or context engineering) refers to methods that improve model behavior by constructing or modifying inputs to an LLM, rather than altering its weights. The current state of the art leverages natural language feedback (shinn2023reflexion; yuksekgonul2024textgrad; agrawal2025gepa). In this paradigm, a language model inspects the current context along with signals such as execution traces, reasoning steps, or validation results, and generates natural language feedback on how the context should be revised. This feedback is then incorporated into the context, enabling iterative adaptation. Representative methods include Reflexion (shinn2023reflexion), which reflects on failures to improve agent planning; TextGrad (yuksekgonul2024textgrad), which optimizes prompts via gradient-like textual feedback; GEPA (agrawal2025gepa), which refines prompts iteratively based on execution traces and achieves strong performance, even surpassing reinforcement learning approaches in some settings; and Dynamic Cheatsheet (krause2019dynamic), which constructs an external memory that accumulates strategies and lessons from past successes and failures during inference. These natural language feedback methods represent a major advance, offering flexible and interpretable signals for improving LLM systems beyond weight updates.
@@ 2.2 Limitations of Existing Context Adaptation Methods
Brevity Bias
A recurring limitation of context adaptation methods is brevity bias: the tendency of optimization to collapse toward short, generic prompts. Gao et al. (gao2025prompt) document this effect in prompt optimization for test generation, where iterative methods repeatedly produced near-identical instructions (e.g.,, “Create unit tests to ensure methods behave as expected”), sacrificing diversity and omitting domain-specific detail. This convergence not only narrows the search space but also propagates recurring errors across iterations, since optimized prompts often inherit the same faults as their seeds. More broadly, such bias undermines performance in domains that demand detailed, context-rich guidance—such as multi-step agents, program synthesis, or knowledge-intensive reasoning—where success hinges on accumulating rather than compressing task-specific insights.
@@
Context Collapse
In a case study on the AppWorld benchmark (trivedi2024appworld), we observe a phenomenon we call context collapse, which arises when an LLM is tasked with fully rewriting the accumulated context at each adaptation step. As the context grows large, the model tends to compress it into much shorter, less informative summaries, causing a dramatic loss of information. For instance, at step 60 the context contained 18,282 tokens and achieved an accuracy of 66.7, but at the very next step it collapsed to just 122 tokens, with accuracy dropping to 57.1—worse than the baseline accuracy of 63.7 without adaptation. While we highlight this through Dynamic Cheatsheet (suzgun2025dynamic), the issue is not specific to that method; rather, it reflects a fundamental risk of end-to-end context rewriting with LLMs, where accumulated knowledge can be abruptly erased instead of preserved.
@@ 3 Agentic Context Engineering (ACE)
We present ACE (Agentic Context Engineering), a framework for scalable and efficient context adaptation in both offline (e.g., system prompt optimization) and online (e.g., test-time memory adaptation) scenarios. Instead of condensing knowledge into terse summaries or static instructions, ACE treats contexts as evolving playbooks that continuously accumulate, refine, and organize strategies over time. Inspired by the agentic design of Dynamic Cheatsheet (suzgun2025dynamic), ACE introduces a structured division of labor across three roles (Figure 4): the Generator, which produces reasoning trajectories; the Reflector, which distills concrete insights from successes and errors; and the Curator, which integrates these insights into structured context updates. This mirrors how humans learn: experimenting, reflecting, and consolidating, while avoiding the bottleneck of overloading a single model with all responsibilities.
@@
To address the limitations of prior methods discussed in §2.2 (notably brevity bias and context collapse) ACE introduces three key innovations: (1) a dedicated Reflector that separates evaluation and insight extraction from curation, improving context quality and downstream performance (§4.6); (2) incremental delta updates3.1) that replace costly monolithic rewrites with localized edits, reducing both latency and compute cost (§4.7); and (3) a grow-and-refine mechanism (§3.2) that balances steady context expansion with redundancy control.
@@ 3.1 Incremental Delta Updates
A core design principle of ACE is to represent context as a collection of structured, itemized bullets, rather than a single monolithic prompt. The concept of a bullet is similar to the concept of a memory entry in LLM memory frameworks like Dynamic Cheatsheet (suzgun2025dynamic) and A-MEM (xu2025mem), but builds on top of that and consists of (1) metadata, including a unique identifier and counters tracking how often it was marked helpful or harmful; and (2) content, capturing a small unit such as a reusable strategy, domain concept, or common failure mode. When solving new problems, the Generator highlights which bullets were useful or misleading, providing feedback that guides the Reflector in proposing corrective updates.
This itemized design enables three properties: (1) localization, so only the relevant bullets are updated; (2) fine-grained retrieval, so the Generator can focus on the most pertinent knowledge; and (3) incremental adaptation, allowing efficient merging, pruning, and de-duplication during inference.
@@ 4 Results
    Enabling High-Performance, Self-Improving Agents. ACE enables agents to self-improve by dynamically refining their input context, both in offline and online settings. It boosts accuracy on the AppWorld benchmark by up to 17.1% by learning to engineer better contexts from execution feedback alone, without needing ground-truth labels. (§4.3)
@@
    Effective by Design. Ablation studies confirm our design choices are key to success, with components like the Reflector, multi-epoch refinement, and incremental delta update each contributing substantial performance gains. (§4.6)
@@ 4.1 Tasks and Datasets
We evaluate ACE on two categories of LLM applications that benefit most from evolving contexts: (1) LLM agent, which require multi-turn reasoning, tool use, and environment interaction; with ACE, agents can accumulate and reuse strategies across episodes and environments; and (2) domain-specific reasoning, which demand mastery of specialized concepts and tactics; we focus on financial analysis as a main case study, and show additional results on medical reasoning and text-to-SQL.
@@
    LLM Agent: AppWorld (trivedi2024appworld) is a suite of autonomous agent tasks involving API understanding, code generation, and environment interaction. It provides a realistic execution environment with common applications and APIs (e.g., email, file system) and tasks of two difficulty levels (normal and challenge). A public leaderboard (AppWorldLeaderboard) tracks performance, where, at the time of submission, the best system achieved only 60.3% average accuracy, highlighting the benchmark’s difficulty and realism.
@@
    Domain-Specific Reasoning: Financial, Medical, and Text-to-SQL Benchmarks We use finance as our main case study in §4.4. For financial analysis, we focus on FiNER (loukas2022finer) and Formula (wang2025finlora), which test LLMs on financial reasoning tasks that rely on the eXtensible Business Reporting Language (XBRL). FiNER requires labeling tokens in XBRL financial documents with one of 139 fine-grained entity types, a key step for financial information extraction in regulated domains. Formula focuses on applying financial concepts and performing computations to answer queries, i.e., numerical reasoning. Beyond finance, we evaluate on two additional domain tasks from StreamBench (wu2024streambench): DDXPlus (fansi2022ddxplus) (medical reasoning) and BIRD-SQL (li2023can) (text-to-SQL).
Evaluation Metrics
For AppWorld, we follow the official benchmark protocol and report Task Goal Completion (TGC) and Scenario Goal Completion (SGC) on both the test-normal and test-challenge splits. For FiNER, Formula and DDXPlus, we follow the original setup and report accuracy, measured as the proportion of predicted answers that exactly match the ground truth. For BIRD-SQL, we use GPT-4o-mini (gpt4omini) under LLM-as-a-judge (zheng2023judging).